### v2

# Diversification and Stochastic Dominance: When All Eggs Are Better Put in One Basket

Abstract
Diversification is widely regarded as a reliable way to reduce risk, yet under certain extreme conditions it can have the opposite effect. A simple and striking example of this phenomenon was recently given by Chen et al. (2025a), who showed that for independent and identically distributed (iid) Pareto risks with infinite mean, any weighted average is larger    in the sense of first-order stochastic dominance    than a single such risk. Our main result    the one-basket theorem  identifies new sufficient conditions under which this reversal occurs for independent but not necessarily identically distributed risks. We compare a weighted average to its corresponding mixture model, which concentrates all exposure on one of the risks, chosen at random. In the iid case, the mixture has the same distribution as any individual risk, thereby recovering the canonical comparison between a diversified portfolio and full exposure to a single component. The theorem enables weight-specific verification of the stochastic dominance relation and yields new applications, including infinite-mean discrete Pareto risks and the St. Petersburg lottery. We further show that these reversals are boundary cases of a broader pattern: diversification always increases the likelihood of exceeding small thresholds, and under specific conditions, this local effect extends globally, resulting in first-order stochastic dominance.
Keywords: diversification, risk pooling, stochastic order, infinite mean
1 Introduction
Over the past century, probabilistic modeling has become a cornerstone of risk management, offering a quantitative framework for assessing and mitigating risk. One of its significant contributions is its formal justification for diversification    the practice of spreading exposure across multiple independent or weakly correlated risks to reduce overall variability. This justification rests on classical results such as the law of large numbers and Modern Portfolio Theory, which demonstrate that diversification can reduce volatility without sacrificing expected returns (Markowitz, 1952). Taken together, these and related developments have come to be seen as confirming a longstanding belief in the benefits of diversification, a notion embedded in common sense and memorably captured by the proverb: “Don’t put all your eggs in one basket.”
1.1 Diversification as a source of risk
Given both this theoretical and intuitive support, it may be surprising that, in certain cases, diversification actually increases the risk. A striking example of this phenomenon was recently established by Chen et al. (2025a), who proved that diversification increases risk in the sense of first-order stochastic dominance when dealing with infinite-mean Pareto risks. More precisely, they showed that for $n\geq 2$ independent and identically distributed (iid) copies $X_{1},\dots,X_{n}$ of a random variable $X$ following a Pareto distribution with shape parameter $\alpha\in(0,1]$, every weighted average of these variables is larger in first-order stochastic dominance than $X$ itself:
$X\leq_{\text{st}}\theta_{1}X_{1}+\dots+\theta_{n}X_{n},$ (1.1)
where $X\leq_{\text{st}}Y$ means $\mathbb{P}(X>x)\leq\mathbb{P}(Y>x)$ for all real $x$.
This result is particularly noteworthy for both the strength of the dominance relation it establishes and the relevance of the distribution it involves. On the one hand, first-order stochastic dominance is arguably the strongest form of stochastic comparison, with broad implications for decision-making. In particular, if the risks represent random losses, then relation (1.1) implies that any decision-maker with an increasing utility function (i.e., someone who favors smaller losses to larger ones) would prefer holding the single risk $X$ rather than the diversified portfolio $\theta_{1}X_{1}+\dots+\theta_{n}X_{n}$  that is, they would choose to put all their eggs in one basket. On the other hand, Pareto distributions with infinite mean are not just theoretical constructs. They arise as useful models for representing certain rare but potentially extreme events, such as the losses from nuclear accidents (Hofert and Wüthrich, 2012; Sornette et al., 2013), cyber and operational risks (Eling and Wirfs, 2019; Eling and Schnell, 2020; Moscadelli, 2004), and fatalities from major earthquakes and pandemics (Clark, 2013; Cirillo and Taleb, 2020).
Although the result of Chen et al. (2025a) is recent and has drawn interest, the phenomenon itself already appeared in earlier work. Embrechts et al. (2002), for instance, proved a case of the stochastic dominance relation for $n=2$ and a Pareto distribution with shape parameter $\alpha=1/2$, and Ibragimov (2005) established it for one-sided stable distributions with infinite mean. Besides this, asymptotic versions of the result have also been established. For example, for risks with regularly varying tails of index $\beta\in(0,1]$, Albrecher et al. (2006) and Embrechts et al. (2009) proved an inequality of the form
$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}>x\right)\geq\mathbb{P}(X>x)$
in the limit as $x\to\infty$.
Returning to the work of Chen et al. (2025a), the authors also extended their result in the same paper to weakly negatively associated and identically distributed (WNAID) super-Pareto risks. In this contribution and those that followed, the prefix super- refers to a class of distributions obtained by applying a convex transformation to the base distribution, often with additional properties such as being increasing and anchored at zero. Subsequent work by Chen and Shneer (2024) extended the stochastic dominance relation to negative lower orthant dependent (NLOD) risks within the so-called $\mathcal{H}$-family. Focusing on the iid case, Müller (2024) and Arab et al. (2024) established the result for super-Cauchy and InvSub risks, respectively.
While each of those results assumed identically distributed risks, Chen et al. (2025c) broadened the scope by relaxing this assumption and characterized diversification through majorization order. Given two positive exposure vectors $(\theta_{1},\dots,\theta_{n})$ and $(\delta_{1},\dots,\delta_{n})$ with equal sum, the former is said to be smaller in majorization order if its components are less dispersed. Using this concept, they showed that the weighted sum of independent but not necessarily identically distributed Pareto risks with infinite mean becomes larger in the sense of first-order stochastic dominance as diversification increases according to the majorization order. Chen et al. (2025b) further extended this result to a broader class of heavy-tailed distributions with infinite mean, showing that this effect occurs beyond the Pareto case.
Chen and Shneer (2024) took a different approach to analyzing diversification with non-identically distributed risks by introducing a framework based on the generalized $r$-mean of the marginal distributions. They established a first-order stochastic dominance result comparing the generalized $r$-mean and the weighted average of NLOD risks with super-Fréchet marginal distributions, which may differ, but all must share a common essential infimum equal to zero. A particularly relevant case occurs at $r=1$, when the generalized $r$-mean of the marginal distributions becomes a mixture model. In that case, their result can be formulated as
$I_{1}X_{1}+\dots+I_{n}X_{n}\leq_{\text{st}}\theta_{1}X_{1}+\dots+\theta_{n}X_{n},$ (1.2)
where $I_{1},\dots,I_{n}$ are Bernoulli random variables such that exactly one of them equal $1$ and the rest are 0, with $\mathbb{P}(I_{i}=1)=\theta_{i}$.
1.2 On the mixture model
The mixture model offers a meaningful generalization of the initial comparison by preserving the idea of concentrating risk in a single position    putting all eggs in one basket    but doing so in a probabilistic manner. This ensures that the mixture remains a relevant benchmark even when the risks are not identically distributed. In the special case of identically distributed risks and the weights sum to $1$, the mixture follows the same distribution as any individual risk, recovering the original result in (1.1) as a particular instance of this broader framework. Moreover, the stochastic dominance relation (1.2) directly implies
$\mathbb{P}\Bigg{(}\sum_{i=1}^{n}\theta_{i}X_{i}>x\Bigg{)}\geq\sum_{i=1}^{n}\theta_{i}\,\mathbb{P}(X_{i}>x)\quad\text{for all }x,$
which relates the distribution of the weighted average    typically lacking a closed-form expression    to the marginal distributions of the risks in a remarkably simple way.
The comparison between a weighted average and the corresponding mixture model also arises in settings that involve randomization. One example appears in the actuarial literature on randomized reinsurance, which differs from standard contracts by introducing exogenous randomness into the coverage mechanism (Albrecher and Cani, 2019; Vincent et al., 2021; Acciaio et al., 2025). Within this framework, the mixture can be viewed as a randomized scheme, where the reinsurer fully covers exactly one loss $X_{i}$, selected at random with probability $\theta_{i}$. The weighted average then corresponds to the deterministic counterpart: a set of quote-share treaties where the reinsurer pays a fixed fraction $\theta_{i}$ of each $X_{i}$. A second example comes from finance, where the weighted average represents a traditional diversified portfolio, while the mixture corresponds to time diversification: in each period, the investor allocates all capital to a single asset, selected at random according to the portfolio weights (Milevsky and Posner, 1996, 1998).
1.3 The content of this paper
The present work builds upon the mixture framework. Our main result is the one-basket theorem, which provides new sufficient conditions for the stochastic dominance relation (1.2) to hold when the risks are independent but not necessarily identically distributed. In particular, our contribution extends the result of Chen and Shneer (2024) in the mixture setting to a broader class of distributions, and removes the requirement that the risks share a common essential infimum.
An important distinction from prior work is that the one-basket theorem does not require the stochastic dominance to hold for all admissible weight vectors. Earlier results identified classes of risks that satisfy relation (1.1) or (1.2) uniformly across all such vectors. In contrast, our approach relaxes this requirement by providing conditions that can be checked for a specific weight vector, even if they are not met for others. This added flexibility allows us to establish stochastic dominance in cases beyond the reach of uniform results. A notable example is the discrete Pareto distribution with infinite mean: although the weighted average of such random variables may fail to dominate a single one under some weight allocations, we show that the equal-weighted average $\frac{1}{n}\sum_{i=1}^{n}X_{i}$ satisfies $X\leq_{\text{st}}\frac{1}{n}\sum_{i=1}^{n}X_{i}$. A similar result holds for the average of St. Petersburg lotteries.
Beyond the main theorem itself, we investigate the types of risks to which it applies. To that end, we introduce the concept of subscalability, which provides a natural interpretation of the key inequalities appearing in the one-basket theorem. This notion serves as the foundation for defining $\theta$-subscalable and completely subscalable risks    two related classes that satisfy the conditions of the theorem. We then analyse the properties of these classes, and relate them to existing families of risks for which dominance relations (1.1) or (1.2) have been previously established.
As a final perspective, we position the one-basket theorem within a broader pattern. Given its sharp contrast with classical diversification principles, the result might initially appear anomalous. However, we show that it arises as the boundary case of a general phenomenon: diversification always increases the likelihood of exceeding small thresholds, and this local effect becomes global when the conditions of the theorem are satisfied, corresponding to first-order stochastic dominance.
The remainder of the paper is organized as follows. Section 2 introduces the setting and notation. Before presenting our main results, Section 3 takes a step back to reaffirm why diversification is typically beneficial in well-behaved settings, using the concept of convex order. Section 4 presents our main results, with particular emphasis on the one-basket theorem. It discusses the relationship of the theorem to previous contributions and illustrates its scope with examples. In Section 5, we introduce the concept of subscalability and use it to characterize classes of risks that satisfy the conditions of the theorem. Section 6 offers an alternative perspective on the theorem, interpreting it as the boundary case of a more general phenomenon. Section 7 concludes. Additional technical details and extended proofs are provided in the appendix.
2 Preliminaries
In what follows, we adopt standard mathematical conventions: “positive” means non-negative, “increasing” means non-decreasing, and comparisons such as “smaller” or “greater” refer to non-strict inequalities. We write $\mathbb{N}=\{1,2,3,\dots\}$ for the set of positive integers.
We consider a probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which all relevant random variables are defined. The random variables of interest are $n\geq 2$ mutually independent random variables $X_{1},\dots,X_{n}$, which we may refer to as risks. Unless stated otherwise, these random variables are assumed to be positive.
Throughout this paper, we work with the survival function of each risk $X_{i}$, defined as $\overline{F}_{i}(x)=\mathbb{P}(X_{i}>x)$. For completeness, we note that the cumulative distribution function is $F_{i}(x)=\mathbb{P}(X_{i}\leq x)$, so that $\overline{F}_{i}(x)=1-F_{i}(x)$. If $X_{i}$ follows distribution $F_{i}$, we may write $X_{i}\sim F_{i}$ or equivalently $X_{i}\sim\overline{F}_{i}$. If two random variables $X$ and $Y$ have the same distribution, we write $X\sim Y$.
We define the weight vector as $\boldsymbol{\theta}:=(\theta_{1},\dots,\theta_{n})$, which is assumed throughout to lie in the open probability simplex:
$\Delta_{n}:=\Bigg{\{}(\theta_{1},\dots,\theta_{n})\in(0,1)^{n}:\sum_{i=1}^{n}\theta_{i}=1\Bigg{\}}.$
The weighted average of the risks is
$\sum_{i=1}^{n}\theta_{i}X_{i}=\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$
It will be referred to as the diversified portfolio.
To define the corresponding mixture model, let $\mathbf{I}=(I_{1},\dots,I_{n})\sim\text{Categorical}(\boldsymbol{\theta})$, independent of $X_{1},\dots,X_{n}$. In other words, a realization of $\mathbf{I}$ has exactly one component equal to $1$ and all others equal to 0, with $\mathbb{P}(I_{i}=1)=\theta_{i}$. The mixture is then
$\sum_{i=1}^{n}I_{i}X_{i}=I_{1}X_{1}+\dots+I_{n}X_{n},$
which selects exactly one of the $X_{i}$ at random according to the weights. In contrast to the diversified portfolio, the mixture concentrates all exposure on a single risk, and will therefore be referred to as the concentrated portfolio.
Although we assume that the weight vector $\boldsymbol{\theta}$ lies in $\Delta_{n}$, meaning that the total weight is exactly one, portfolios with total weight strictly less than one can still be represented. This can be done by setting one or more of the risks $X_{i}$ to be trivial  that is, almost surely zero.
The survival function of the concentrated portfolio is given by
$\mathbb{P}\left(\sum_{i=1}^{n}I_{i}X_{i}>x\right)=\sum_{i=1}^{n}\theta_{i}\,\mathbb{P}(X_{i}>x)=\sum_{i=1}^{n}\theta_{i}\,\overline{F}_{i}(x).$
In the special case where the risks are identically distributed with common survival function $\overline{F}(x)$, this simplifies to
$\mathbb{P}\left(\sum_{i=1}^{n}I_{i}X_{i}>x\right)=\sum_{i=1}^{n}\theta_{i}\,\overline{F}(x)=\overline{F}(x),$
since the weights sum to one whenever $\boldsymbol{\theta}\in\Delta_{n}$. Thus, when the risks are identically distributed, the concentrated portfolio has the same distribution as any individual risk.
Let $n]:=\{1,\dots,n\}$. For any nonempty subset $\mu\subseteq[n]$, we define its subset weight as
$\theta_{\mu}:=\sum_{i\in\mu}\theta_{i}.$
With this notation, a single subscript (e.g., $\theta_{i}$) refers to an individual weight, whereas a subscript corresponding to a nonempty subset (e.g., $\theta_{\mu}$) denotes the total weight over that subset.
We now recall the definition of first-order stochastic dominance used throughout the paper. Given two random variables $X$ and $Y$, we say that $X$ is smaller than $Y$ in the sense of first-order stochastic dominance, written $X\leq_{\text{st}}Y$, if
$\mathbb{P}(X>x)\leq\mathbb{P}(Y>x)\quad\text{for all }x\in\mathbb{R}.$
Since all random variables considered in this paper are non-negative, the inequality above holds trivially on the interval $(-\infty,0)$, where both survival functions equal 1. Accordingly, whenever we establish first-order stochastic dominance, we will do so by verifying the inequality on the domain $0,\infty)$. The only exception is in Example 5.3, where a transformed variable may take on negative values. In that case, dominance is established by applying a lemma that does not require checking the inequality directly.
We conclude this section by recalling the following two basic properties of first-order stochastic dominance, which will be used throughout.
Lemma 2.1.
(Closure under increasing transformations) If $X\leq_{\text{st}}Y$ and $f$ is any increasing function, then $f(X)\leq_{\text{st}}f(Y)$.
(Closure under convolution) Let $X_{1},\dots,X_{m}$ and $Y_{1},\dots,Y_{m}$ be two sets of independent random variables, and suppose that $X_{i}\leq_{\text{st}}Y_{i}$ for each $i=1,\dots,m$. Then $X_{1}+\dots+X_{m}\leq_{\text{st}}Y_{1}+\dots+Y_{m}$.
For proofs, as well as a comprehensive treatment of stochastic orders, see the classic textbooks by Müller and Stoyan (2002) and Shaked and Shanthikumar (2007).
3 The classical view on diversification
In the introduction, we reviewed several results from the literature challenging the usual intuition about diversification, showing that under specific conditions, it can actually increase risk. Before turning to our main contribution, which further extends these findings, we first take a step back to reaffirm the view on why diversification is typically beneficial. In particular, we formalize this intuition within our setting using the concept of convex order, which offers a natural framework to compare the diversified and concentrated portfolios.
3.1 A reminder on convex order
Convex order is a well-established method for comparing random variable and has numerous applications in risk analysis and decision theory. Formally, for two random variables $X$ and $Y$, we say that $X$ is smaller than $Y$ in convex order, denoted by $X\leq_{\text{cx}}Y$, if
$\mathbb{E}[c(X)]\leq\mathbb{E}[c(Y)]$
holds for all convex functions $c$ for which the expectations are well-defined.
Convex order is usually considered in contexts where expectations are finite, in which case it can be interpreted as a way to compare the variability of random quantities that share the same expectation. This interpretation arises from two key observations. First, non-constant convex functions emphasize extreme values, so the inequality in the definition suggests that $Y$ tends to take on more extreme realizations than $X$. Second, since both functions $c(x)=x$ and c(x)=xc(x)=-x are convex, it follows that if $X\leq_{\text{cx}}Y$ and the expectations exist, then $\mathbb{E}[X]=\mathbb{E}[Y]$. When, in addition, these expectations are finite, they provide a meaningful common reference point. Taken together, these observations confirm that convex order, in the finite-mean case, captures differences in variability around a shared expectation.
This notion of risk contrasts with that captured by first-order stochastic dominance, which compares the relative location of distributions. Specifically, $X\leq_{\text{st}}Y$ means that $X$ is less likely than $Y$ to exceed any given threshold, reflecting a shift in probability mass toward smaller values. Roughly speaking, while first-order stochastic dominance captures a difference in size, convex order reflects (under finite expectations) differences in variability between random variables of the same size.
Several important consequences follow from convex order, again assuming the relevant expectations are finite. For example, a risk-averse decision-maker will prefer the loss $X$ over $Y$ whenever $X\leq_{\text{cx}}Y$, since risk aversion corresponds to a utility function that is convex over the loss domain. Likewise, since the function $c(x)=(x-a)^{2}$ is convex for any fixed $a\in\mathbb{R}$, the convex order relation implies that $X$ has a smaller variance than $Y$. This is closely related to the classical justification for diversification provided by the law of large numbers and Modern Portfolio Theory, as discussed in the introduction.
Further applications of convex order, along with a detailed treatment, can be found in (Müller and Stoyan, 2002) and (Shaked and Shanthikumar, 2007).
3.2 Convex order in our setting
To show that a convex order relation holds in our setting, the key observation is that the diversified portfolio can be expressed as the conditional expectation of the concentrated one, given the random variables $\mathbf{X}:=(X_{1},\dots,X_{n})$. That is, letting $P_{C}:=\sum_{i=1}^{n}I_{i}X_{i}$ and $P_{D}:=\sum_{i=1}^{n}\theta_{i}X_{i}$, we have
$\mathbb{E}[P_{C}|\mathbf{X}]=P_{D}.$
Applying Jensen’s inequality to any convex function $c$ for which expectations are well-defined, we then obtain
$\mathbb{E}[c(P_{D})]=\mathbb{E}[c(\mathbb{E}[P_{C}|\mathbf{X}])]\leq\mathbb{E}[\mathbb{E}[c(P_{C})|\mathbf{X}]]=\mathbb{E}[c(P_{C})],$
which establishes the convex order relation.
This argument yields the following lemma, a standard consequence of Jensen’s inequality.
Lemma 3.1.
Consider $n\geq 2$ random variables $X_{1},\dots,X_{n}$. Given a weight vector $\boldsymbol{\theta}\in\Delta_{n}$, let $\mathbf{I}\sim\text{Categorical}(\boldsymbol{\theta})$, independent of $X_{1},\dots,X_{n}$. Then the following convex order relation holds:
$\theta_{1}X_{1}+\cdots+\theta_{n}X_{n}\leq_{\text{cx}}I_{1}X_{1}+\cdots+I_{n}X_{n}.$
Hence, under finite expectations, the interpretation and consequences of convex order outlined in the previous subsection apply. In particular, the diversified portfolio exhibits less variability than the concentrated one while preserving the same mean, confirming the classical intuition that diversification reduces risk in well-behaved settings.
By contrast, when expectations are infinite, they no longer provide meaningful information about the location of risks, and the interpretation above breaks down (see (Côté and Wang, 2025) for a detailed analysis of convex order under non-finite expectations). In these situations, counterintuitive phenomena can emerge, including cases where the diversified portfolio is actually larger than the concentrated one in first-order stochastic dominance, and thus presents more risk. The next section introduces the one-basket theorem, which provides sufficient conditions under which this surprising reversal occurs.
4 Main results
In this section, we present our core findings. We begin with the single-risk case, which paves the way for the multi-risk framework. We then establish a general inequality comparing the distribution of the diversified portfolio to that of the individual risks. This inequality serves as an intermediate step leading to the one-basket theorem, which provides sufficient conditions for the diversified portfolio to be larger than the concentrated one in the sense of first-order stochastic dominance. Finally, we illustrate the result with examples and highlight connections to prior work.
4.1 The case of a single risk
Consider the simple case where only one of the risks is non-trivial, and the others are almost surely zero. For ease of notation, we temporarily omit subscripts. The mixture then reduces to $IX$, and the weighted average simplifies to $\theta X$, where $I\sim\text{Bernoulli}(\theta)$ is independent of $X$, and $\theta\in(0,1)$. In this setting, $IX$ can be interpreted as a lottery that delivers $X$ with probability $\theta$ and 0 otherwise, while $\theta X$ scales down the risk deterministically.
To determine when first-order stochastic dominance holds between these two random variables, we compare their respective survival functions. The probability that $IX$ exceeds a threshold $x\geq 0$ is $\mathbb{P}(IX>x)=\theta\,\overline{F}(x)$, while for $\theta X$, it is $\mathbb{P}(\theta X>x)=\overline{F}(x/\theta)$. By the definition of first-order stochastic dominance, this leads to the following result:
Lemma 4.1.
Let $X$ be a positive random variable. Given $\theta\in(0,1)$, let $I\sim\text{Bernoulli}(\theta)$ be independent of $X$. Then $IX\leq_{\text{st}}\theta X$ if and only if the inequality $\theta\,\overline{F}(x)\leq\overline{F}(x/\theta)$ holds for all $x\geq 0$.
The inequality $\theta\overline{F}(x)\leq\overline{F}(x/\theta)$ thus fully determines when $IX$ is smaller than $\theta X$ in the sense of first-order stochastic dominance, and hints at a pattern that may extend to the multi-risk setting. In fact, it will be central to both the proof and the formulation of the next theorem, which requires keeping track of the points where this inequality holds, for each risk and across various subset weights. To that end, we define for each $X_{i}$ the region
$r_{i}(\theta):=\{x\geq 0:\theta\,\overline{F}_{i}(x)\leq\overline{F}_{i}(x/\theta)\},\qquad\theta\in(0,1).$
In Section 5, we examine in more detail the class of risks whose survival functions satisfy this inequality pointwise for all $x\geq 0$, for some $\theta\in(0,1)$. Among other properties, we will show that if such a $\theta$ exists for a non-trivial risk, then that risk must have infinite mean.
4.2 The general case
When multiple risks are involved, extending the single-risk result requires a more refined approach. A careful partition of the sample space (see Lemma C.2 in the appendix) allows us to bound the survival function of the diversified portfolio from below over the region
$\mathcal{R}(\boldsymbol{\theta}):=\bigcap_{i=1}^{n}\bigcap_{\{i\}\subseteq\mu\subset[n]}r_{i}(\theta_{\mu}),\qquad\boldsymbol{\theta}\in\Delta_{n},$
where $\theta_{\mu}=\sum_{i\in\mu}\theta_{i}$ is the subset weight of $\mu$, as previously defined.
Over that region, the lower bound takes the form of a weighted average of the marginal survival functions, as established in the following theorem.
Theorem 4.1.
Consider $n\geq 2$ independent positive random variables $X_{1},\dots,X_{n}$ with survival functions $\overline{F}_{1},\dots,\overline{F}_{n}$. Given a weight vector $\boldsymbol{\theta}\in\Delta_{n}$, the inequality
$\mathbb{P}\Bigg{(}\sum_{i=1}^{n}\theta_{i}X_{i}>x\Bigg{)}\geq\sum_{i=1}^{n}\theta_{i}\,\mathbb{P}(X_{i}>x)$
holds for all $x\in\mathcal{R}(\boldsymbol{\theta})$.
The proof is given in Appendix C.
That theorem has several key implications. Notably, as we show in Section 6.2, the region $\mathcal{R}(\boldsymbol{\theta})$ always includes a non-trivial interval $0,t(\boldsymbol{\theta}))$, which ensures that the inequality holds for small values of $x$ in all cases. However, our main focus here is on the situations where the inequality extends to all positive values.
Recall from Section 2 that the survival function of the concentrated portfolio equals the weighted average of the marginal survival functions. Therefore, Theorem 4.1 implies that the inequality
$\mathbb{P}\left(\sum_{i=1}^{n}I_{i}X_{i}>x\right)\leq\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}>x\right)$
holds for all $x\geq 0$ whenever $\mathcal{R}(\boldsymbol{\theta})=[0,\infty)$, so the diversified portfolio is larger than the concentrated one in the sense of first-order stochastic dominance. By the definition of $\mathcal{R}(\boldsymbol{\theta})$, this occurs precisely when, for each $i\in[n]$ and every $\mu\subset[n]$ with $\mu\supseteq\{i\}$, the inequality $\theta_{\mu}\,\overline{F}_{i}(x)\leq\overline{F}_{i}(x/\theta_{\mu})$ holds for all $x\geq 0$.
This leads to our main result:
Theorem 4.2 (One-basket theorem).
Consider $n\geq 2$ independent positive random variables $X_{1},\dots,X_{n}$ with survival functions $\overline{F}_{1},\dots,\overline{F}_{n}$. Given a weight vector $\boldsymbol{\theta}\in\Delta_{n}$, let $\mathbf{I}\sim\text{Categorical}(\boldsymbol{\theta})$, independent of $X_{1},\dots,X_{n}$. Suppose that for each $i\in[n]$ and every $\mu\subset[n]$ with $\mu\supseteq\{i\}$, the condition
$\theta_{\mu}\,\overline{F}_{i}(x)\leq\overline{F}_{i}(x/\theta_{\mu})\qquad\text{for all }x\geq 0,$ (4.1)
is satisfied. Then the following first-order stochastic dominance relation holds:
$I_{1}X_{1}+\dots+I_{n}X_{n}\leq_{\text{st}}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$ (4.2)
This theorem shows that, under suitable conditions, the diversified portfolio carries more risk than the concentrated one in the sense of first-order stochastic dominance. In such cases, when the risks $X_{1},\dots,X_{n}$ represent losses or other adverse outcomes, any risk-averse agent would prefer holding the concentrated portfolio, which places all exposure on a single risk selected at random    in other words, they would prefer to have all their eggs put in one basket. Hence the name of the theorem.
4.3 Discussion
The one-basket theorem connects to earlier results in the literature on stochastic dominance and risk aggregation. We now explore these connections and illustrate the scope of the theorem through two examples. Additional examples follow in the next section.
As already mentioned in the introduction, the mixture case (corresponding to the concentrated portfolio) was also studied by Chen and Shneer (2024), who established relation (4.2) for NLOD (negative lower orthant dependent) super-Fréchet risks with possibly different marginal distributions, under the assumption that all risks share a common essential infimum equal to zero. In contrast, our result assumes independence, but applies to a broader class of distributions (see Proposition 5.7), and does not require a common lower bound.
The following example gives a concrete case where the result applies even though the risks have different essential infima.
Example 4.1 (Non-identically distributed Pareto risks).
Let $X_{1},\dots,X_{n}$ be $n\geq 2$ independent random variables such that each $X_{i}$ follows a Pareto distribution with shape parameter $\alpha_{i}\in(0,1]$ and scale parameter $\rho_{i}>0$. The survival function of $X_{i}$ is given by $\overline{F}_{i}(x)=(\rho_{i}/x)^{\alpha_{i}}$ for $x\geq\rho_{i}$ and $\overline{F}_{i}(x)=1$ otherwise. Since $\alpha_{i}\leq 1$, each $X_{i}$ has infinite mean. The essential infimum of each risk is $\rho_{i}$, which may vary across $i$.
It can be verified that for each $i$, condition (4.1) holds for all $\theta\in(0,1)$. Since for any weight vector $\boldsymbol{\theta}\in\Delta_{n}$ the resulting subset weights $\theta_{\mu}$ all lie in $(0,1)$, it follows that the condition is satisfied for all relevant $i$ and $\mu$. As a result, the one-basket theorem applies in this setting, and the diversified portfolio induced by any $\boldsymbol{\theta}\in\Delta_{n}$ is larger than the corresponding concentrated one in the sense of first-order stochastic dominance, as given by relation (4.2). This yields a particularly simple lower bound for the survival function of the diversified portfolio:
$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}>x\right)\geq\sum_{i=1}^{n}\gamma_{i}(x)\,x^{-\alpha_{i}},$
where $\gamma_{i}(x):=\mathbf{1}_{x\geq\rho_{i}}\,\theta_{i}\,\rho_{i}^{\alpha_{i}}$. $\blacksquare$
While the previous example highlights the flexibility of the theorem, our result also covers the traditional setting where the risks are identically distributed. In this scenario, the concentrated portfolio has the same distribution as any individual risk, and the comparison reduces to evaluating whether a single risk is stochastically dominated by a weighted average of its iid copies. As mentioned earlier, this case has been studied by Ibragimov (2005), Chen et al. (2025a), Chen and Shneer (2024), Müller (2024), and Arab et al. (2024).
A notable departure from prior work is that the one-basket theorem does not require the stochastic dominance relation to hold uniformly across all admissible weight vectors. Instead, it identifies a set of conditions that can be tested for any specific weight vector. If these conditions are satisfied, first-order stochastic dominance is guaranteed, regardless of whether they fail under other allocations. This relaxation allows us to establish dominance results for distributions beyond the reach of existing results.
To illustrate this feature, we consider a discrete version of the Pareto distribution, for which the conditions of the theorem hold for some weight vectors but fails for others.
Example 4.2 (Average of iid discrete Pareto risks).
Let $X_{1},\dots,X_{n}$ be iid copies of a random variable $X$ with survival function given by $\overline{F}(x)=(\lfloor x\rfloor+2)^{-1}$ for $x\geq 0$ and $\overline{F}(x)=1$ otherwise, where $\lfloor x\rfloor$ denotes the greatest integer less than or equal to $x$. The distribution of $X$ coincides with that of $\lfloor\widetilde{Y}\rfloor-1$, where $\widetilde{Y}$ follows Pareto distribution with shape and scale parameters equal to $1$. Thus, $X$ is a discrete Pareto random variable, shifted to have support starting at 0.
In the identically distributed setting, relation (4.2) simplifies to $X\leq_{\text{st}}\sum_{i=1}^{n}\theta_{i}X_{i}$. As shown in Lemma A.1 (appendix), the inequality $\theta\,\overline{F}(x)\leq\overline{F}(x/\theta)$ holds for all $x\geq 0$ if and only if the weight $\theta\in(0,1)$ lies in the set $\mathcal{A}:=\left(0,\tfrac{1}{2}\right]\cup\left\{\tfrac{k+1}{2k+1}:k\in\mathbb{N}\right\}$. Since $\mathcal{A}$ is a strict subset of $(0,1)$, the one-basket theorem fails to apply under certain allocations, and first-order stochastic dominance is not guaranteed in general. For example, if $n=2$ and $\theta_{1}=0.9$, then $\mathbb{P}(X>\theta_{1})=0.5>\frac{61}{132}=\mathbb{P}(\theta_{1}X_{1}+\theta_{2}X2>\theta_{1})$, which is not compatible with $X\leq_{\text{st}}\theta_{1}X_{1}+\theta_{2}X_{2}$.
However, in the case of equal weights, the stochastic dominance relation does hold. Let $\theta_{i}=\frac{1}{n}$, so that the diversified portfolio is the average of the risks $\overline{X}_{n}:=\frac{1}{n}\sum_{i=1}^{n}X_{i}$, and relation (4.2) becomes $X\leq_{\text{st}}\overline{X}_{n}$. The relevant subset weights are then $\{\frac{i}{n}:i=1,\dots,n-1\}$, which all lie in $\mathcal{A}$ when $n=2$ or $n=3$, and the theorem yields $X\leq_{\text{st}}\overline{X}_{2}$ and $X\leq_{\text{st}}\overline{X}_{3}$. For $n\geq 4$, at least one subset weight (namely $\frac{n-1}{n}$) falls outside $\mathcal{A}$, so the theorem no longer applies directly. Nonetheless, an inductive argument presented in Proposition A.1 establishes that the stochastic dominance relation $X\leq_{\text{st}}\overline{X}_{n}$ continues to hold for all $n\geq 2$. The same stochastic dominance relation holds for the unshifted discrete Pareto random variable $Y=\lfloor\widetilde{Y}\rfloor$, as shown later in Example 5.3. $\blacksquare$
Finally, our setting reveals a noteworthy implication for randomized reinsurance, as studied by Albrecher and Cani (2019); Vincent et al. (2021); Acciaio et al. (2025), and described in Section 1.2. For illustration, it is sufficient to consider the single-risk case. Let $X\sim\overline{F}$ represent an insurance loss. Under a classical quota-share treaty, the insurer cedes a fixed fraction $\theta X$ to the reinsurer and retains $(1-\theta)X$. In contrast, consider a randomized reinsurance scheme in which the reinsurer covers the full loss $X$ with probability $\theta$, and pays nothing otherwise. The ceded and retained losses are then $IX$ and (1I)X(1-I)X, where $I\sim\text{Bernoulli}(\theta)$ is independent of $X$. In line with the result established in Lemma 4.1, if the survival function $\overline{F}$ satisfies the inequalities
$\theta\,\overline{F}(x)\leq\overline{F}(x/\theta)\quad\text{and}\quad(1-\theta)\,\overline{F}(x)\leq\overline{F}(x/(1-\theta))\quad\text{for all }x\geq 0,$
then both $IX\leq_{\text{st}}\theta X$ and $(1-I)X\leq_{\text{st}}(1-\theta)X$ hold. In that case, both the reinsurer and the insurer are better off under the randomized scheme in the sense of first-order stochastic dominance    a perhaps surprising outcome, since adding exogenous randomness leads to better risk-sharing.
Although we focused on the single-risk case for clarity, similar improvements may arise in multivariate settings, as shown by the one-basket theorem. Moreover, extensions to other randomized contract designs    beyond the all-or-nothing coverage    may also prove beneficial under suitable conditions, offering interesting directions for future research.
5 Subscalability
In the previous section, we showed that a first-order stochastic dominance relation holds between the diversified and concentrated portfolios when certain conditions on survival functions are met. Here, we explore the types of risks that satisfy these conditions by introducing the concept of subscalability, which underpins the definitions $\theta$-subscalable and completely subscalable risks    two nested classes covered by the one-basket theorem. We then examine their properties and relate them to risk families studied in earlier work.
5.1 The concept of subscalability
The one-basket theorem relies on a key condition: that the survival functions of the risks satisfy inequalities of the form $\theta\,\overline{F}(x)\leq\overline{F}(x/\theta)$, where $\theta\in(0,1)$. Intuitively, if the survival function of a risk $X$ satisfies this inequality at a given point $x$, it means that scaling down the risk by a factor $\theta$ reduces the probability of exceeding $x$ less than proportionally. That is,
$\theta\,\overline{F}(x)\leq\overline{F}(x/\theta)\quad\Longleftrightarrow\quad\theta\,\mathbb{P}(X>x)\leq\mathbb{P}(\theta X>x).$
At such points, we may say that the risk (or its survival function) resists scaling    a property that we refer to as subscalability.
5.2 $\theta$-subscalable risks
Building on this concept, we now define the class of $\theta$-subscalable risks.
Definition 5.1.
For a given $\theta\in(0,1)$, a positive random variable $X\sim\overline{F}$ is said to be $\theta$-subscalable if the inequality $\theta\,\overline{F}(x)\leq\overline{F}(x/\theta)$ holds for all $x\geq 0$. We also refer to the corresponding distribution as $\theta$-subscalable.
This class comprises a variety of distributions, including both discrete and continuous cases. Some distributions are $\theta$-subscalable for all scaling factors $\theta\in(0,1)$. This includes the Pareto distribution with infinite mean, as established in Example 4.1. Distributions with this stronger property are studied further in Section 5.3. In contrast, other distributions are $\theta$-subscalable only for some values of $\theta$, and not for others. It is the case for the discrete Pareto distribution discussed in Example 4.2, which was shown to be $\theta$-subscalable if and only if $\theta\in\mathcal{A}$.
A natural question then arises regarding the set of $\theta$ for which a given distribution is $\theta$-subscalable: how small can it be? While this set can, of course, be empty for some distributions, the following lemma shows that if it contains at least one value, then it must also contain an infinite sequence of smaller ones.
Lemma 5.1.
Let $\theta\in(0,1)$ and suppose that the random variable $X$ is $\theta$-subscalable. Then $X$ is also $\theta^{k}$-subscalable, for all $k\in\mathbb{N}$.
Proof.
Since the inequality $\theta\,\overline{F}(x)\leq\overline{F}(x/\theta)$ holds for all $x\geq 0$, applying it at $x/\theta^{j}$, for each $j=0,\dots,k-1$, gives
$\theta\,\overline{F}(x/\theta^{j})\leq\overline{F}(x/\theta^{j+1})\quad\text{for all }x\geq 0.$
Composing these inequalities yields
$\theta^{k}\,\overline{F}(x)\leq\overline{F}(x/\theta^{k})\quad\text{for all }x\geq 0,$
which proves that $X$ is $\theta^{k}$-subscalable. ∎
Thus, subscalability at a single $\theta\in(0,1)$ entails subscalability at all smaller scaling factors of the form $\theta^{k}$ with $k\in\mathbb{N}$, making the geometric sequence $\{\theta^{k}:k\in\mathbb{N}\}$ the minimal set for which subscalability must hold.
Interestingly, for some distributions, that minimal set is also complete: the distribution is subscalable exactly at these values, and at no others. We illustrate this in the next example, which also presents an opportunity to apply the one-basket theorem and establish a new first-order stochastic dominance result.
Example 5.1 (St. Petersburg lottery).
The St. Petersburg lottery is a classical construct in probability theory that yields a distribution with infinite mean. In this game, a fair coin is tossed until the first head appears, and the player then receives a payoff of $X=2^{m}$, where $m$ denotes the number of tosses. The resulting distribution has a probability mass function $\mathbb{P}(X=2^{m})=2^{-m}$, $m\in\mathbb{N}$, leading to $\mathbb{E}[X]=\sum_{m=1}^{\infty}2^{m}\,2^{-m}=\infty$. The survival function of $X$ is given by $\overline{F}(x)=2^{-\lfloor\log_{2}x\rfloor}$ for $x\geq 2$ and $\overline{F}(x)=1$ otherwise.
As established in Lemma B.1 (appendix), $X$ is $\theta$-subscalable if and only if $\theta$ belongs to $\mathcal{B}:=\{2^{-k}:k\in\mathbb{N}\}$. Thus, the St. Petersburg lottery provides an instance of a distribution for which a geometric set of scaling factors is not only minimal, but also complete.
This example also yields a direct application of the one-basket theorem. Indeed, we show in Proposition B.1 that for iid St. Petersburg lotteries $X_{1},\dots,X_{n}$, the average $\overline{X}_{n}=\frac{1}{n}\sum_{i=1}^{n}X_{i}$ satisfies the stochastic dominance relation $X\leq_{\text{st}}\overline{X}_{n}$ for all $n\in\{2^{k}:k\in\mathbb{N}\}$. $\blacksquare$
We now establish a second basic property of the class, which links $\theta$-subscalability to the tail behavior of the distribution.
Lemma 5.2.
Let $\theta\in(0,1)$, and suppose $X$ is a non-trivial $\theta$-subscalable random variable. Then $\mathbb{E}[X]=\infty$.
Proof.
Since $X$ is non-trivial, there exists a point $x>0$ such that $x\,\overline{F}(x)>0$. Moreover, as $X$ is $\theta$-subscalable, Lemma 5.1 yields
$\theta^{k}\,\overline{F}(x)\leq\overline{F}(x/\theta^{k})\quad\Longrightarrow\quad x\,\overline{F}(x)\leq x/\theta^{k}\,\overline{F}(x/\theta^{k}),$
for all $k\in\mathbb{N}$. Thus $t\,\overline{F}(t)$ remains strictly positive as $t\to\infty$, which implies $\mathbb{E}[X]=\infty$, since $X$ is positive. ∎
With this lemma, we can relate $\theta$-subscalability to heavy-tailedness, which plays a central role in the modeling of extreme risks. A distribution is said to be heavy-tailed if its survival function decays more slowly than any exponential    that is, if $\lim_{x\rightarrow\infty}\,e^{tx}\,\overline{F}(x)=\infty$ for all $t>0$. Since this property is implied by infinite mean, Lemma 5.2 shows that every non-trivial $\theta$-subscalable risk is necessarily heavy-tailed. For background on this concept, see (Resnick, 2007) or (Embrechts et al., 2013).
To conclude this subsection, we mention that not all random variables with infinite mean are $\theta$-subscalable. For example, the survival function $\overline{F}(x)=\mathbf{1}_{\{x<e\}}+\mathbf{1}_{\{x\geq e\}}\,\frac{1}{x\ln x}$ yields a positive random variable with infinite mean, yet it fails to satisfy the inequality defining $\theta$-subscalability for any $\theta\in(0,1)$.
5.3 Completely subscalable risks
A stronger condition than $\theta$-subscalability arises when the inequality $\theta\,\overline{F}(x)\leq\overline{F}(x/\theta)$ holds for all $x\geq 0$ and all $\theta\in(0,1)$. We define such risks as completely subscalable.
Definition 5.2.
A positive random variable $X$ is said to be completely subscalable if the inequality $\theta\overline{F}(x)\leq\overline{F}(x/\theta)$ holds for all $x\geq 0$ and all $\theta\in(0,1)$. We also refer to the corresponding distribution as completely subscalable.
It follows directly from the definition that a completely subscalable risk is $\theta$-subscalable for all $\theta\in(0,1)$. By Lemma 5.2, this implies that any non-trivial completely subscalable risk must have infinite mean.
An equivalent characterization of complete subscalability is given by the following lemma.
Lemma 5.3.
A positive random variable $X\sim\overline{F}$ is completely subscalable if and only if the function $h(x):=x\,\overline{F}(x)$ is increasing on $(0,\infty)$.
Proof.
Assume first that $X$ is completely subscalable. Then for any $0<x<y$, set $\theta:=x/y\in(0,1)$. The defining inequality gives $x\,\overline{F}(x)\leq y\,\overline{F}(y)$, which shows that $h$ is increasing on $(0,\infty)$. Conversely, suppose that $h$ is increasing on $(0,\infty)$. Fix $x>0$ and let $y:=x/\theta$. Then by monotonicity of $h$, we have $\theta\,\overline{F}(x)\leq\overline{F}(x/\theta)$. Since this inequality also holds at $x=0$, we conclude that $X$ is completely subscalable. ∎
As a consequence, we obtain the following continuity property.
Lemma 5.4.
Let $X\sim\overline{F}$ be completely subscalable. Then $\overline{F}$ is continuous on $(0,\infty)$.
Proof.
If $\overline{F}$ had a downward jump at any $x>0$, then $h(x)=x\,\overline{F}(x)$ would jump down as well, contradicting the monotonicity property established in Lemma 5.3. Hence $\overline{F}$ must be continuous on $(0,\infty)$. ∎
The next lemma establishes a useful closure property.
Lemma 5.5.
Let $X\sim\overline{F}$ be completely subscalable and let $g$ be an increasing convex function with $g(0)=0$. Then the random variable $Y:=g(X)$ is also completely subscalable.
Proof.
Let $\overline{G}(x)=\mathbb{P}(Y>x)=\mathbb{P}(g(X)>x)$. By convexity and anchoring at zero, we have $\theta g(X)\geq g(\theta X)$, so
$\overline{G}(x/\theta)=\mathbb{P}(\theta g(X)>x)\geq\mathbb{P}(g(\theta X)>x)=\overline{F}(g^{-1}(x)/\theta).$
By complete subscalability of $X$, it follows that
$\overline{F}(g^{-1}(x)/\theta)\geq\theta\,\overline{F}(g^{-1}(x))=\theta\,\overline{G}(x).$
Thus $\theta\,\overline{G}(x)\leq\overline{G}(x/\theta)$, proving that $Y$ is completely subscalable. ∎
We now relate the class of completely subscalable risks to the super-Fréchet family introduced by Chen and Shneer (2024), under which their stochastic dominance result was derived. As mentioned earlier, the authors proved a broader stochastic dominance result involving the generalized $r$-mean and the weighted average of NLOD super-Fréchet risks with possibly different distributions but common essential infimum of 0. A special case of this result, corresponding to $r=1$, yields the first-order stochastic dominance relation:
$I_{1}X_{1}+\cdots+I_{n}X_{n}\leq_{\text{st}}\theta_{1}X_{1}+\cdots+\theta_{n}X_{n},$
as in the one-basket theorem.
Definition 5.3.
[…]
For convenience, we denote the classes of super-Fréchet and completely subscalable risks by $\mathcal{S}_{\mathcal{F}}$ and $\mathcal{CS}$, respectively.
The next lemma shows that the Fréchet(1)(1) distribution belongs to $\mathcal{CS}$.
Lemma 5.6.
[…]
Proof.
[…]
$\theta(1-e^{-1/x})\leq 1-e^{-\theta/x}\quad\Longrightarrow\quad\theta\,\overline{F}(x)\leq\overline{F}(x/\theta),$
which verifies the inequality defining complete subscalability for all $x>0$. Given that $\overline{F}(0)=1$, the inequality also holds at $x=0$. Hence the Fréchet(1)(1) distribution is completely subscalable. ∎
This yields the following relationship between the two classes:
Lemma 5.7.
$\mathcal{S}_{F}\subset\mathcal{CS}$.
Proof.
[…]
Many classical heavy-tailed distributions belong to $\mathcal{S}_{\mathcal{F}}$. For example, the Burr, log-logistic, Lomax and paralogistic distributions with infinite mean are shown in Example 4 of Chen and Shneer (2024) to be super-Fréchet. It follows from Lemma 5.7 that they are also completely subscalable.
5.4 On the iid case
As previously noted, when the risks are identically distributed, the concentrated portfolio has the same distribution as any of the marginal risks, and the dominance relation from the one-basket theorem simplifies to
$X\leq_{\text{st}}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$ (5.1)
In the iid setting, if a random variable $X$ satisfies this relation, then so does any increasing convex transformation of $X$, say $f(X)$. That property follows from the preservation of first-order stochastic dominance under increasing transformations (Lemma 2.1), together with Jensen’s inequality:
$f(X)\leq_{\text{st}}f(\theta_{1}X_{1}+\dots+\theta_{n}X_{n})\leq\theta_{1}f(X_{1})+\dots+\theta_{n}f(X_{n}).$
The implication above is a classical consequence of Jensen’s inequality, which we state here for a fixed weight vector, in line with the setting of the one-basket theorem:
Lemma 5.8.
If relation (5.1) holds for some weight vector $\boldsymbol{\theta}\in\Delta_{n}$, then we also have
$f(X)\leq_{\text{st}}\theta_{1}f(X_{1})+\dots+\theta_{n}f(X_{n}),$
for any increasing convex function $f$.
This lemma offers two immediate benefits. First, it allows us to derive new cases from known ones    for example, if the Pareto distribution with infinite mean satisfies the inequality, then so does any increasing convex transformation of it, without requiring a separate proof. Second, it extends the result to new settings, such as distributions with negative support. The following examples illustrate both points.
Example 5.2.
As established in (Chen et al., 2025a), for any weight vector $\boldsymbol{\theta}\in\Delta_{n}$, the stochastic dominance relation $X\leq_{\text{st}}\sum_{i=1}^{n}\theta_{i}X_{i}$ holds when $X$ follows a Pareto distribution with shape parameter in $(0,1]$, so that it has infinite mean. This also follows from Example 4.1, by setting all shape parameters $\alpha_{i}\in(0,1]$ and all scale parameters $\rho_{i}>0$ to be equal. By the lemma above, the same relation holds for any increasing convex transformation of such a Pareto variable. For example, since the exponential function is increasing and convex, the dominance relation remains valid for the log-Pareto variable $Y=e^{X}$. $\blacksquare$
Example 5.3.
In Example 4.2, we showed that the stochastic dominance relation $X\leq_{\text{st}}\overline{X}_{n}$ holds for all $n\geq 2$ when $X$ has a survival function given $\overline{F}=(\lfloor x\rfloor+2)^{-1}$, for $x\geq 0$ and $\overline{F}(x)=1$ otherwise. By applying the lemma, this relation also holds for any increasing convex transformation $Y=f(X)$. For instance, since the function $f(x)=x+a$ is increasing convex for any constant $a\in\mathbb{R}$, the shifted variable $Y=X+a$ also satisfies the dominance relation. When $a<0$ the support of $Y$ includes at least one strictly negative value. In the case $a=1$, the survival function of $Y$ is given by $\overline{F}(x)=(\lfloor x\rfloor+1)^{-1}$ for $x\geq 1$, and $\overline{F}(x)=1$ otherwise. This coincides with the survival function of the discrete random variable $\lfloor\widetilde{Y}\rfloor$, where $\widetilde{Y}$ follows Pareto distribution with shape and scale parameters equal to $1$. $\blacksquare$
As noted earlier, the iid case has received particular attention in previous works, under the stronger requirement that the stochastic dominance relation holds for all weights $\theta_{1},...,\theta_{n}\geq 0$ with $\sum_{i=1}^{n}\theta_{i}=1$. Ibragimov (2005) established this result for one-sided stable distributions with infinite mean. Chen et al. (2025a) later proved it for super-Pareto risks, while Chen and Shneer (2024) extended it to the so-called $\mathcal{H}$-family. Lastly, Müller (2024) and Arab et al. (2024) showed it for super-Cauchy and InvSub risks, respectively.
To date, the super-Cauchy and InvSub families represent the largest known explicit classes for which the stochastic dominance relation holds uniformly across all admissible weight vectors. The two classes overlap but are not nested, as illustrated by examples in Section 4 of (Müller, 2024). A super-Cauchy risk is a random variable that can be written as $c(Y)$, where $Y$, where $Y$ is a standard Cauchy, with survival function $\overline{F}(x)=\frac{1}{2}-\frac{1}{\pi}\arctan(x)$, $x\in\mathbb{R}$, and $c$ is a convex function. Importantly, $c$ is not required to be increasing, anchored at zero, or even positive. As a result, the class includes random variables with support beyond the positive reals    such as the Cauchy distribution itself.
Regarding the InvSub class, a preliminary comparison suggests that, as currently defined, it does not strictly contain the completely subscalable risks, though it could with a slight generalization of the definition. However, some clarification is needed, as Arab et al. (2024) present two distinct formulations of the InvSub class that are not equivalent, leaving the precise scope of the class somewhat ambiguous at this stage.
6 The one-basket theorem as a boundary case
The one-basket theorem identifies sufficient conditions under which diversification increases risk in the sense of first-order stochastic dominance. This result contradicts classical intuition and may appear anomalous at first glance. However, in this section, we show that rather than being a pathological outlier, it arises as the boundary case of a general phenomenon: diversification always increases risk locally near the origin, and this effect becomes global when the conditions of the one-basket theorem are met.
We proceed in three steps. First, we prove that every positive risk exhibits subscalability over a non-trivial interval near the origin. Next, we demonstrate that subscalability in this region gives rise to what can be seen as a local version of the one-basket theorem, valid in all cases near zero. To measure how far this effect is guaranteed, we introduce the threshold function $t(\boldsymbol{\theta})$. Finally, we consider the case where $t(\boldsymbol{\theta})$ is infinite: in this setting, the effect extends globally, yielding the one-basket theorem.
6.1 Subscalability near the origin
We introduced the notion of subscalability in Section 5.1, to describe the situation in which scaling down a risk $X\sim\overline{F}$ by a factor $\theta\in(0,1)$ reduces the probability of exceeding a given threshold $x$ less than proportionally:
$\theta\,\mathbb{P}(X>x)\leq\mathbb{P}(\theta X>x)\quad\Longleftrightarrow\quad\theta\,\overline{F}(x)\leq\overline{F}(x/\theta).$
In the one-basket theorem, this inequality is required to hold globally    that is, for all $x\geq 0$, across all risks and all relevant subset weights.
Such a global requirement is strong and fails in many cases. However, we now show that the inequality always holds locally near the origin, for any individual risk and any $\theta\in(0,1)$.
This simply follows from the right-continuity of survival functions: the inequality is trivially satisfied at $x=0$, and by right-continuity of $\overline{F}$, it continues to hold on some interval $0,\varepsilon)$ with $\varepsilon>0$.
To formalize it, we define for each risk $X_{i}$ and each $\theta\in(0,1)$:
$t_{i}(\theta):=\sup\left\{t\geq 0:\theta\,\overline{F}_{i}(x)\leq\overline{F}_{i}(x/\theta)\right\}.$
Then we have:
Lemma 6.1.
Let $X_{i}\sim\overline{F}_{i}$ be a positive random variable. Then $t_{i}(\theta)>0$ for any $\theta\in(0,1)$.
6.2 A local version of the one-basket theorem
While the one-basket theorem establishes first-order stochastic dominance under specific conditions, we now show that a local version of this result always holds: the diversified portfolio is more likely to exceed small thresholds than the concentrated one, regardless of whether the conditions of the theorem are satisfied.
To see this, recall that by Theorem 4.1, given any weight vector $\boldsymbol{\theta}\in\Delta_{n}$, the inequality
$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}>x\right)\geq\sum_{i=1}^{n}\theta_{i}\,\mathbb{P}(X_{i}>x)$ (6.1)
holds for all $x$ in the region $\mathcal{R}(\boldsymbol{\theta})$, defined as
$\mathcal{R}(\boldsymbol{\theta})=\bigcap_{i=1}^{n}\bigcap_{\{i\}\subseteq\mu\subset[n]}r_{i}(\theta_{\mu}),$
with
$r_{i}(\theta)=\left\{x\geq 0:\overline{F}_{i}(x/\theta)\geq\theta\,\overline{F}_{i}(x)\right\}.$
By construction, each region $r_{i}(\theta_{\mu})$ contains the interval $0,t_{i}(\theta_{\mu}))$. Therefore,
$\mathcal{R}(\boldsymbol{\theta})\supseteq[0,t(\boldsymbol{\theta})),\quad\text{where}\quad t(\boldsymbol{\theta}):=\min_{i=1}^{n}\min_{\{i\}\subseteq\mu\subset[n]}t_{i}(\theta_{\mu}).$
Moreover, since each $t_{i}(\theta_{\mu})$ is strictly positive by Lemma 6.1, we have
$t(\boldsymbol{\theta})>0.$
Thus, inequality (6.1) holds for all $x\in[0,t(\boldsymbol{\theta}))$, and since the right-hand side equals the survival function of the concentrated portfolio,
$\mathbb{P}\left(\sum_{i=1}^{n}I_{i}X_{i}>x\right)=\sum_{i=1}^{n}\theta_{i}\,\mathbb{P}(X_{i}>x),$
we obtain the following proposition:
Proposition 6.1.
Consider $n\geq 2$ independent positive random variables $X_{1},\dots,X_{n}$ with survival functions $\overline{F}_{1},\dots,\overline{F}_{n}$. Given a weight vector $\boldsymbol{\theta}\in\Delta_{n}$, the region $\mathcal{R}(\boldsymbol{\theta})$ contains the interval $0,t(\boldsymbol{\theta}))$, where $t(\boldsymbol{\theta})>0$. Therefore, the inequality
$\mathbb{P}\left(\sum_{i=1}^{n}I_{i}X_{i}>x\right)\leq\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}>x\right)$ (6.2)
holds for all $x\in[0,t(\boldsymbol{\theta}))$.
In line with the initial claim, this result establishes a local version of the one-basket theorem. Indeed, recall that the theorem asserts the stochastic dominance relation
$I_{1}X_{1}+\cdots+I_{n}X_{n}\leq_{\text{st}}\theta_{1}X_{1}+\cdots+\theta_{n}X_{n},$
which, by definition, is equivalent to the inequality (6.2) holding for all $x\geq 0$. In other words, when the theorem applies, the diversified portfolio is more likely than the concentrated one to exceed any given threshold. Proposition 6.1 shows that this behavior is always guaranteed locally: the inequality holds on a non-trivial interval $0,t(\boldsymbol{\theta}))$, meaning that the diversified portfolio is more likely to exceed sufficiently small thresholds    even when global stochastic dominance fails.
6.3 When the local effect becomes global
As we just established, a local version of the one-basket theorem always holds near the origin. This is no coincidence: the two phenomena are directly connected. In fact, the one-basket theorem characterizes exactly the setting in which that local effect extends globally.
This connection becomes clear upon examining Proposition 6.1 under the assumption that $t(\boldsymbol{\theta})=\infty$. Then, the proposition guarantees that inequality (6.2) holds for all $x\geq 0$, yielding first-order stochastic dominance:
$I_{1}X_{1}+\dots+I_{n}X_{n}\leq_{\text{st}}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$
By definition, the threshold $t(\boldsymbol{\theta})$ is infinite if and only if the condition
$\theta_{\mu}\,\overline{F}_{i}(x)\leq\overline{F}_{i}(x/\theta_{\mu})\qquad\text{for all }x\geq 0$
is satisfied for each $i\in[n]$ and for every $\mu\subset[n]$ with $\mu\subseteq\{i\}$. These are exactly the conditions and the conclusion of the one-basket theorem.
The theorem thus appears as a special instance of the proposition, confirming the connection between the local and global phenomena. Seen through this lens, the theorem emerges not as an anomaly, but as the boundary case of a broader effect: diversification always increases the likelihood of exceeding small thresholds, and while this is typically confined to a limited region near the origin, under specific conditions it extends globally, giving rise to first-order stochastic dominance.
7 Conclusion
This paper provides new sufficient conditions under which diversification increases risk in the sense of first-order stochastic dominance. Our main result    the one-basket theorem    extends previous work by accommodating non-identically distributed risks, and by allowing the dominance relation to be established for specific weight vectors, rather than requiring it to hold uniformly across all admissible allocations. This flexibility is especially useful in discrete settings, where uniform results usually fail. In particular, it allowed us to establish new stochastic dominance results for discrete Pareto risks and for St. Petersburg lotteries.
To analyse and interpret the conditions of the theorem, we introduced the concept of subscalability, along with the related notions of $\theta$-subscalable and completely subscalable risks. These classes offer intuitive and tractable criteria for identifying risks that exhibit the type of behavior described by the one-basket theorem, and they provide a useful structure for studying their properties. In addition, they help connect our findings to results previously established in the literature.
Finally, we showed that the one-basket theorem arises as the boundary case of a broader phenomenon: diversification always increases the likelihood of exceeding small thresholds, regardless of distributional assumptions. Under certain conditions, this local effect extends globally, resulting in first-order stochastic dominance. This perspective situates the theorem within a larger pattern and clarifies how diversification can fail.
Appendix
Appendix A The iid discrete Pareto example
This section supports the analysis in Example 4.2, which concerns a discrete Pareto distribution with survival function
$\overline{F}(x)=\begin{cases}1,&x<0,\\ (\lfloor x\rfloor+2)^{-1},&x\geq 0.\end{cases}$ (A.1)
We first identify the set of weights leading to $\theta$-subscalability.
Lemma A.1.
Let $\theta\in(0,1)$ and set $\overline{F}$ as in (A.1). Then the inequality $\theta\,\overline{F}(x)\leq\overline{F}(x/\theta)$ holds for all $x\geq 0$ if and only if $\theta\in\mathcal{A}$, where
$\mathcal{A}=\left(0,\tfrac{1}{2}\right]\cup\left\{\tfrac{k+1}{2k+1}:k\in\mathbb{N}\right\}.$
Proof.
First, observe that at $x=0$, the inequality becomes $\theta\overline{F}(0)\leq\overline{F}(0)$, which trivially holds for all $\theta\in(0,1)$. We therefore focus on the case $x>0$.
Let $x>0$ and define
$\ell:=\lfloor x\rfloor+1,\qquad m:=\left\lfloor x/\theta\right\rfloor+1,$
so that the inequality rewrites as
$\theta(m+1)\leq\ell+1.$ (A.2)
To analyse this, fix $m\in\mathbb{N}$ and consider the range of $x$ such that $m=\lfloor x/\theta\rfloor+1$, i.e.,
$x\in[(m-1)\theta,\,m\theta).$ (A.3)
Within this interval, $\ell=\lfloor x\rfloor+1$ attains its minimum at $x=(m-1)\theta$, namely
$\ell=\lfloor(m-1)\theta\rfloor+1.$ (A.4)
Since increasing $\ell$ makes inequality (A.2) easier to satisfy, it suffices to verify it at this minimal value. The inequality becomes
$\theta(m+1)\leq\lfloor(m-1)\theta\rfloor+2\quad\Longleftrightarrow\quad\operatorname{frac}((m-1)\theta)\leq 2(1-\theta),$
where $\operatorname{frac}(z)=z-\lfloor z\rfloor$ denotes the fractional part. Thus, the original inequality holds for all $x>0$ if and only if
$\operatorname{frac}((m-1)\theta)\leq 2(1-\theta)\qquad\text{for all }m\in\mathbb{N}.$ (A.5)
We now split into cases according to the value of $\theta$.
Case 1: $\theta\in(0,\tfrac{1}{2}]$. In this range, $2(1-\theta)\geq 1$, while $\operatorname{frac}(z)<1$ for all $z\in\mathbb{R}$, so condition (A.5) is always satisfied.
Case 2: $\theta\in(\tfrac{1}{2},1)$ irrational. By Kronecker’s theorem (see, e.g., Theorem 12.2.2 in (Miller and Takloo-Bighash, 2006)), the set $\{\operatorname{frac}((m-1)\theta):m\in\mathbb{N}\}$ is dense in $(0,1)$. Since $2(1-\theta)<1$, the inequality (A.5) fails for some $m$.
Case 3: $\theta\in(\tfrac{1}{2},1)$ rational. Write $\theta=\frac{a}{b}$, where $a,b\in\mathbb{N}$ are coprime and satisfy $2a>b$ and $a<b$. For each $m\in\mathbb{N}$, we have
$(m-1)\theta=\frac{(m-1)a}{b},\quad\text{so}\quad\operatorname{frac}((m-1)\theta)=\frac{r}{b},$
where $r=(m-1)a\bmod b$. Since $a$ and $b$ are coprime, the values of $r$ range over $\{0,1,\dots,b-1\}$ as $m$ varies, making the maximum fractional part (b1)/b(b-1)/b. Therefore, (A.5) requires
$\frac{b-1}{b}\leq 2\left(1-\frac{a}{b}\right)\quad\Longleftrightarrow\quad 2a\leq b+1.$
Combined with $2a>b$, this implies $2a=b+1$. Letting $b=2k+1$ for $k\in\mathbb{N}$ gives $a=k+1$, so
$\theta=\frac{k+1}{2k+1}.$
Putting these cases together yields the result. ∎
With this, we are now ready to show that the relation $X\leq_{\mathrm{st}}\overline{X}_{n}$ holds for all $n\geq 2$.
Proposition A.1.
Let $X$ be a random variable with survival function $\overline{F}$ as in (A.1) and let $\overline{X}_{n}=\frac{1}{n}\sum_{i=1}^{n}X_{i}$ be the average of $n\geq 2$ iid copies of $X$. Then
$X\leq_{\mathrm{st}}\overline{X}_{n}\quad\text{for all }n\geq 2.$
Proof.
We apply the one-basket theorem to the identically distributed case. In this setting, the relevant subset weights are given by $\{\frac{i}{n}:i=1,\dots,n-1\}$, and condition
$\theta\,\overline{F}(x)\leq\overline{F}(x/\theta)\quad\text{for all }x\geq 0,$
must hold for each of them. According to Lemma A.1, this condition holds if and only if the weight $\theta\in(0,1)$ lies in $\mathcal{A}=(0,\tfrac{1}{2}]\cup\{\tfrac{k+1}{2k+1}:k\in\mathbb{N}\}$.
When $n=2$, the only relevant subset weight is $\frac{1}{2}\in\mathcal{A}$, and when $n=3$, the weights are $\frac{1}{3},\frac{2}{3}\in\mathcal{A}$. Therefore, the one-basket theorem applies and yields
$X\leq_{\mathrm{st}}\overline{X}_{2}\qquad\text{and}\qquad X\leq_{\mathrm{st}}\overline{X}_{3}.$ (A.6)
Now fix $n\geq 4$, and assume as the induction hypothesis that the inequality holds for all integers $j=2,\dots,n-1$. Let $m:=\lfloor n/2\rfloor$, and define
$Y_{1}:=\frac{1}{m}\sum_{i=1}^{m}X_{i},\qquad Y_{2}:=\frac{1}{n-m}\sum_{i=m+1}^{n}X_{i}.$
Then $Y_{1}\sim\overline{X}_{m}$ and $Y_{2}\sim\overline{X}_{n-m}$. Since $n\geq 4$, we have $2\leq m\leq n-2$, so both $m$ and nmn-m are at least $2$. This ensures that the induction hypothesis applies to both $Y_{1}$ and $Y_{2}$, and so we have
$X\leq_{\mathrm{st}}Y_{1}\qquad\text{and}\qquad X\leq_{\mathrm{st}}Y_{2}.$
Define $\theta_{1}:=\frac{m}{n}$, $\theta_{2}:=1-\theta_{1}=\frac{n-m}{n}$. By stochastic dominance being preserved under positive scaling and convolution (Lemma 2.1), we obtain
$\theta_{1}X_{1}^{\prime}+\theta_{2}X_{2}^{\prime}\leq_{\mathrm{st}}\theta_{1}Y_{1}+\theta_{2}Y_{2}=\overline{X}_{n},$
where $X_{1}^{\prime}$ and $X_{2}^{\prime}$ are independent copies of $X$.
To apply the one-basket theorem to the left-hand side, we verify that the relevant weights (namely $\theta_{1}$ and $\theta_{2}$) lie in $\mathcal{A}$. If $n$ is even, then $\theta_{1}=\theta_{2}=\tfrac{1}{2}$. If $n$ is odd, we can write $n=2k+1$ for some $k\in\mathbb{N}$, which gives $\theta_{1}=\frac{k}{2k+1}$ and $\theta_{2}=\frac{k+1}{2k+1}$. In all cases, the weights belong to $\mathcal{A}$, so the one-basket theorem applies and gives
$X\leq_{\mathrm{st}}\theta_{1}X_{1}^{\prime}+\theta_{2}X_{2}^{\prime}.$
Combining with the earlier inequality by transitivity, we conclude
$X\leq_{\mathrm{st}}\overline{X}_{n},$
which completes the induction step.
Together with the base cases $n=2$ and $n=3$, this establishes the result for all $n\geq 2$.
Appendix B The St. Petersburg lottery example
This section provides formal results and proofs related to the subscalability and stochastic dominance properties of the St. Petersburg lottery from Example 5.1. The associated survival function is given by
$\overline{F}(x)=\begin{cases}1,&x<2,\\ 2^{-\lfloor\log_{2}x\rfloor},&x\geq 2.\end{cases}$ (B.1)
We first characterize the set of values $\theta$ for which $\theta$-subscalability holds.
Lemma B.1.
Let $\theta\in(0,1)$ and set $\overline{F}$ as in (B.1). Then the inequality $\theta\,\overline{F}(x)\leq\overline{F}(x/\theta)$ holds for all $x\geq 0$ if and only if $\theta\in\mathcal{B}$, where
$\mathcal{B}=\{2^{-k}:k\in\mathbb{N}\}.$
Proof.
First, suppose that $\theta\in\mathcal{B}$, so that $\theta=2^{-k}$ for some $k\in\mathbb{N}$. A case-by-case check using the piecewise definition of $\overline{F}$ confirms that the inequality $\theta\,\overline{F}(x)\leq\overline{F}(x/\theta)$ holds for all $x\geq 0$.
Conversely, suppose $\theta\in(0,1)\setminus\mathcal{B}$. Then $\theta=2^{-(k-1+\delta)}$ for some $k\in\mathbb{N}$ and $0<\delta<1$. In this case, the inequality fails, for instance at $x=2^{1-\delta}$, since
$\theta\,\overline{F}(x)=2^{-(k-1+\delta)}>2^{-k}=\overline{F}(x/\theta).$
Hence, the inequality holds if and only if $\theta\in\mathcal{B}$. ∎
With this, we now establish the following first-order stochastic dominance relations for iid St. Petersburg lotteries:
Proposition B.1.
Let $X$ be a random variable with survival function $\overline{F}$ as in (B.1). Let further $\overline{X}_{n}=\frac{1}{n}\sum_{i=1}^{n}X_{i}$ be the average of $n\geq 2$ iid copies of $X$. Then
$X\leq_{\mathrm{st}}\overline{X}_{n}\quad\text{for all }n\in\{2^{k}:k\in\mathbb{N}\}.$
Proof.
We begin with the base case $n=2$. The one-basket theorem applies if the condition
$\theta\,\overline{F}(x)\leq\overline{F}(x/\theta)\quad\text{for all }x\geq 0,$
is satisfied for the weight $\theta=\frac{1}{2}$. According to Lemma B.1, this condition is satisfied. The one-basket theorem thus applies and we obtain
$X\leq_{\text{st}}\overline{X}_{2}.$
For the remaining cases, we proceed by induction on $k$. Suppose the result holds for some $k\in\mathbb{N}$, that is
$X\leq_{\text{st}}\overline{X}_{2^{k}}.$
We aim to prove that this yields
$X\leq_{\text{st}}\overline{X}_{2^{k+1}}.$
Let $X_{1},\dots,X_{2^{k+1}}$ be iid copies of $X$. Group them into two independent blocks of size $2^{k}$, as
$Y_{1}:=\frac{1}{2^{k}}\sum_{i=1}^{2^{k}}X_{i},\qquad Y_{2}:=\frac{1}{2^{k}}\sum_{i=2^{k}+1}^{2^{k+1}}X_{i}.$
Then $Y_{1}$ and $Y_{2}$ are iid and both distributed as $\overline{X}_{2^{k}}$. By the induction hypothesis,
$X\leq_{\text{st}}Y_{1}\quad\text{and}\quad X\leq_{\text{st}}Y_{2}.$
Let $X_{1}^{\prime}$ and $X_{2}^{\prime}$ be independent copies of $X$. Since first-order stochastic dominance is preserved under positive scaling and convolution (Lemma 2.1), we obtain
$\frac{1}{2}(X_{1}^{\prime}+X_{2}^{\prime})\leq_{\text{st}}\frac{1}{2}(Y_{1}+Y_{2}).$
Observe that $\frac{1}{2}(X_{1}^{\prime}+X_{2}^{\prime})\sim\overline{X}_{2}$, and $\frac{1}{2}(Y_{1}+Y_{2})\sim\overline{X}_{2^{k+1}}$, so the above relation implies
$\overline{X}_{2}\leq_{\text{st}}\overline{X}_{2^{k+1}}.$
Now, applying the base case $X\leq_{\text{st}}\overline{X}_{2}$ and transitivity, we conclude that
$X\leq_{\text{st}}\overline{X}_{2^{k+1}},$
which completes the induction step.
Thus, by induction, the stochastic dominance relation $X\leq_{\text{st}}\overline{X}_{n}$ holds for all $n=2^{k}$ with $k\in\mathbb{N}$. ∎
Appendix C Proof of Theorem 4.1
This appendix provides the full proof of Theorem 1, restated below for convenience.
Theorem 4.1 (Restated).
Consider $n\geq 2$ independent positive random variables $X_{1},\dots,X_{n}$ with survival functions $\overline{F}_{1},\dots,\overline{F}_{n}$. Given a weight vector $\boldsymbol{\theta}\in\Delta_{n}$, the inequality
$\mathbb{P}\Bigg{(}\sum_{i=1}^{n}\theta_{i}X_{i}>x\Bigg{)}\geq\sum_{i=1}^{n}\theta_{i}\,\mathbb{P}(X_{i}>x)$
holds for all $x\in\mathcal{R}(\boldsymbol{\theta})$.
We begin by introducing notation and establishing two preparatory lemmas that will be used in the proof.
C.1 Preliminaries
Let $\mathcal{S}\subseteq[n]$ be a non-empty index subset. For each $\mu\subseteq\mathcal{S}$ define
$A_{\mu}(u):=\bigcap_{i\in\mu}\{X_{i}>u\},\quad u\in\mathbb{R},$
the event that all variables with indices in $\mu$ exceed the threshold $u$.
Next, let
$\mathbf{u}:=(u_{\lambda}:\emptyset\subset\lambda\subseteq\mathcal{S})$
be a vector of real thresholds indexed by the nonempty subsets of $\mathcal{S}$. For each such vector and subset $\mu\subseteq\mathcal{S}$, define the event
$B_{\mu}(\mathbf{u}):=A_{\mu}(u_{\mu})\cap\bigcap_{\mu\subset\lambda\subseteq\mathcal{S}}\overline{A_{\lambda}(u_{\lambda})}.$
Thus $B_{\mu}(\mathbf{u})$ occurs precisely when the variables indexed by $\mu$ exceed the threshold $u_{\mu}$, while no strict superset $\lambda\supset\mu$ within $\mathcal{S}$ exceeds its corresponding threshold $u_{\lambda}$.
The following two properties of $A_{\mu}(u)$ will be used throughout:
  • Decomposability: $A_{\mu}(u)=A_{\sigma}(u)\cap A_{\mu\setminus\sigma}(u)$, for all $\sigma\subseteq\mu\subseteq\mathcal{S}$.
  • Monotonicity: $A_{\mu}(u_{1})\subseteq A_{\mu}(u_{2})$, for all $u_{1}\geq u_{2}$.
C.2 Auxiliary lemmas
The first lemma provides an alternative representation of the event $B_{\mu}(\mathbf{u})$ under a monotonicity condition on the thresholds, which will be useful in later arguments.
Lemma C.1.
Let $\emptyset\subset\mu\subseteq\mathcal{S}\subseteq[n]$, and suppose the thresholds satisfy
$u_{\mu}\geq u_{\lambda}\qquad\text{for all }\mu\subset\lambda\subseteq\mathcal{S}.$
Then for every $\sigma\subseteq\mu$, the event $B_{\mu}(\mathbf{u})$ can be rewritten as
$B_{\mu}(\mathbf{u})=A_{\mu}(u_{\mu})\cap\bigcap_{\mu\subset\lambda\subseteq\mathcal{S}}\overline{A_{\lambda\setminus\sigma}(u_{\lambda})}.$
Proof.
Write $A_{\lambda}(u_{\lambda})=A_{\sigma}\cap A_{\lambda\setminus\sigma}(u_{\lambda})$. By De Morgan, we have
$\overline{A_{\lambda}(u_{\lambda})}=\overline{A_{\sigma}(u_{\lambda})}\cup\overline{A_{\lambda\setminus\sigma}(u_{\lambda})},$
which in $B_{\mu}(\mathbf{u})$ leads to
$B_{\mu}(\mathbf{u})=A_{\mu}(u_{\mu})\cap\bigcap_{\mu\subset\lambda\subseteq\mathcal{S}}\left(\overline{A_{\sigma}(u_{\lambda})}\cup\overline{A_{\lambda\setminus\sigma}(u_{\lambda})}\right)$
Because $A_{\mu}(u_{\mu})\subseteq A_{\sigma}(u_{\mu})\subseteq A_{\sigma}(u_{\lambda})$, the intersection $A_{\mu}(u_{\mu})\cap\overline{A_{\sigma}(u_{\lambda})}$ is empty, leaving only the stated expression. ∎
The next lemma establishes that the events $B_{\mu}(\mathbf{u})$ partition the sample space under the same threshold condition. This result is central to applying the law of total probability in the main argument.
Lemma C.2.
Let $\mathcal{S}\subseteq[n]$, and suppose the thresholds satisfy
$u_{\mu}\geq u_{\lambda}\qquad\text{for all }\mu\subset\lambda\subseteq\mathcal{S}.$
Then the collection $\{B_{\mu}(\mathbf{u}):\mu\subseteq\mathcal{S}\}$ forms a partition of the sample space $\Omega$.
Proof.
For distinct $\mu_{1},\mu_{2}\subseteq\mathcal{S}$, put $\tau:=\mu_{1}\cup\mu_{2}$, so that $\mu_{1},\mu_{2}\subset\tau$. Lemma C.1 yields
$B_{\mu_{1}}(\mathbf{u})=A_{\mu_{1}}(u_{\mu_{1}})\cap\bigcap_{\mu_{1}\subset\lambda\subseteq\mathcal{S}}\overline{A_{\lambda\setminus\mu_{1}}(u_{\lambda})}\subseteq\overline{A_{\tau\setminus\mu_{1}}(u_{\tau})}=\overline{A_{\mu_{2}}(u_{\tau})},$
Using the definition of $B_{\mu}(\mathbf{u})$, we can then write
$B_{\mu_{2}}(\mathbf{u})\subseteq A_{\mu_{2}}(u_{\mu_{2}})\subseteq A_{\mu_{2}}(u_{\tau}),$
where the last inclusion uses $u_{\mu_{2}}\geq u_{\tau}$ and the monotonicity property of $A_{\mu}(u)$. Hence the events are pairwise disjoint, and there remains to show that their union covers $\Omega$.
For each outcome $\omega\in\bigcup_{\emptyset\subset\mu\subseteq\mathcal{S}}A_{\mu}(u_{\mu})$ define
$\mu^{*}(\omega):=\bigcup_{\begin{subarray}{c}\emptyset\subset\mu\subseteq\mathcal{S}\\ \omega\in A_{\mu}(u_{\mu})\end{subarray}}\mu.$
Then $\omega\in A_{\mu^{*}}(u_{\mu^{*}})$ and $\omega\notin A_{\lambda}(u_{\lambda})$ for all $\lambda\supset\mu^{*}$, so $\omega\in B_{\mu^{*}}(\mathbf{u})$. Therefore
$\bigcup_{\emptyset\subset\mu\subseteq\mathcal{S}}B_{\mu}(\mathbf{u})\supseteq\bigcup_{\emptyset\subset\mu\subseteq\mathcal{S}}A_{\mu}(u_{\mu}).$
The reverse inclusion is immediate because $B_{\mu}(\mathbf{u})\subseteq A_{\mu}(u_{\mu})$, leading to
$\bigcup_{\emptyset\subset\mu\subseteq\mathcal{S}}B_{\mu}(\mathbf{u})=\bigcup_{\emptyset\subset\mu\subseteq\mathcal{S}}A_{\mu}(u_{\mu}).$
Using de Morgan,
$B_{\emptyset}(\mathbf{u})=\overline{\bigcup_{\emptyset\subset\mu\subseteq\mathcal{S}}A_{\mu}(u_{\mu})}.$
and so adding $B_{\emptyset}(\mathbf{u})$ to the union gives
$\bigcup_{\mu\subseteq\mathcal{S}}B_{\mu}(\mathbf{u})=\Omega.$
C.3 Main argument
We now prove Theorem 4.1 using the preparatory material above. The argument proceeds by applying the law of total probability, exploiting the partition structure, and using independence.
Fix $x\in\mathcal{R}(\boldsymbol{\theta})$. Define
$u_{\mu}:=x/\theta_{\mu},$
and let
$\mathbf{u}=(u_{\mu}:\emptyset\subset\mu\subseteq[n])$
be the resulting threshold vector, as in the definition of $B_{\mu}(\mathbf{u})$. Since $\boldsymbol{\theta}\in\Delta_{n}$, each subset weight satisfies $\theta_{\mu}>0$, ensuring the thresholds are well-defined. Because $\theta_{\mu}\leq\theta_{\lambda}$ for all $\emptyset\subset\mu\subset\lambda\subseteq[n]$, and as $x\in\mathcal{R}(\boldsymbol{\theta})$ implies $x\geq 0$, it follows that
$u_{\mu}\geq u_{\lambda}.$
Consequently, Lemma C.2 applies, and the events $\{B_{\mu}(\mathbf{u}):\mu\subseteq[n]\}$ partition the sample space $\Omega$.
Applying the law of total probability, we obtain
$\mathbb{P}\Bigg{(}\sum_{i=1}^{n}\theta_{i}X_{i}>x\Bigg{)}=\sum_{\mu\subseteq[n]}\mathbb{P}\Bigg{(}\sum_{i=1}^{n}\theta_{i}X_{i}>x,B_{\mu}(\mathbf{u})\Bigg{)},$
and dropping the term corresponding to $\mu=\emptyset$ thus yields the inequality
$\mathbb{P}\Bigg{(}\sum_{i=1}^{n}\theta_{i}X_{i}>x\Bigg{)}\geq\sum_{\emptyset\subset\mu\subseteq[n]}\mathbb{P}\Bigg{(}\sum_{i=1}^{n}\theta_{i}X_{i}>x,B_{\mu}(\mathbf{u})\Bigg{)}.$
Because each $\theta_{i}$ and each $X_{i}$ are positive, the weighted sum $\sum_{i\in\mu}\theta_{i}X_{i}$ can only increase when we add terms. Together with the given definitions, this results in the chain of inclusions
$B_{\mu}(\mathbf{u})\subseteq A_{\mu}(u_{\mu})\subseteq\Bigg{\{}\sum_{i\in\mu}\theta_{i}X_{i}>x\Bigg{\}}\subseteq\Bigg{\{}\sum_{i=1}^{n}\theta_{i}X_{i}>x\Bigg{\}}.$
Therefore $B_{\mu}(\mathbf{u})\subseteq\{\sum_{i=1}^{n}\theta_{i}X_{i}>x\}$, and we have
$\mathbb{P}\Bigg{(}\sum_{i=1}^{n}\theta_{i}X_{i}>x,B_{\mu}(\mathbf{u})\Bigg{)}=\mathbb{P}(B_{\mu}(\mathbf{u})),$
which simplifies the inequality to
$\mathbb{P}\Bigg{(}\sum_{i=1}^{n}\theta_{i}X_{i}>x\Bigg{)}\geq\sum_{\emptyset\subset\mu\subseteq[n]}\mathbb{P}(B_{\mu}(\mathbf{u})).$
Noting $\sum_{i\in\mu}\frac{\theta_{i}}{\theta_{\mu}}=1$ for each nonempty $\mu$, we may insert this identity into each summand:
$\mathbb{P}\Bigg{(}\sum_{i=1}^{n}\theta_{i}X_{i}>x\Bigg{)}\geq\sum_{\emptyset\subset\mu\subseteq[n]}\sum_{i\in\mu}\frac{\theta_{i}}{\theta_{\mu}}\,\mathbb{P}(B_{\mu}(\mathbf{u})).$
After changing the order of summation and appropriately redefining the indices, we obtain
$\mathbb{P}\Bigg{(}\sum_{i=1}^{n}\theta_{i}X_{i}>x\Bigg{)}\geq\sum_{i=1}^{n}\sum_{\mu\subseteq[n]\setminus\{i\}}\frac{\theta_{i}}{\theta_{\mu}}\,\mathbb{P}(B_{\mu\cup\{i\}}(\mathbf{u})).$
Now, define
$v_{\mu}^{(i)}:=u_{\mu\cup\{i\}},$
and let
$\mathbf{v}^{(i)}:=(v_{\mu}^{(i)}:\emptyset\subset\mu\subseteq[n]\setminus\{i\})$
be the resulting new threshold vector. With this, we can write
$B_{\mu\cup\{i\}}(\mathbf{u})=\{X_{i}>x/\theta_{\mu\cup\{i\}}\}\cap B_{\mu}(\mathbf{v}^{(i)}).$
Since $X_{i}$ and $X_{j}$ are independent for any $i\neq j$ and $B_{\mu}(\mathbf{v}^{(i)})$ does not involve $X_{i}$ for all $\mu\subseteq[n]\setminus\{i\}$, the events $\{X_{i}>x/\theta_{\mu\cup\{i\}}\}$ and $B_{\mu}(\mathbf{v}^{(i)})$ are independent. Hence,
$\mathbb{P}(B_{\mu\cup\{i\}}(\mathbf{u}))=\overline{F}_{i}(x/\theta_{\mu\cup\{i\}})\,\mathbb{P}(B_{\mu}(\mathbf{v}^{(i)})).$
As $x\in\mathcal{R}(\boldsymbol{\theta})$, this simplifies to
$\mathbb{P}(B_{\mu\cup\{i\}}(\mathbf{u}))=\theta_{\mu\cup\{i\}}\,\overline{F}_{i}(x)\,\mathbb{P}(B_{\mu}(\mathbf{v}^{(i)})).$
By injecting this back into the inequality, we obtain
$\mathbb{P}\Bigg{(}\sum_{i=1}^{n}\theta_{i}X_{i}>x\Bigg{)}\geq\sum_{i=1}^{n}\theta_{i}\,\overline{F}_{i}(x)\sum_{\mu\subseteq[n]\setminus\{i\}}\mathbb{P}(B_{\mu}(\mathbf{v}^{(i)}))$
From the properties of the thresholds in $\mathbf{u}$ that we established earlier, it follows that the elements of $\mathbf{v}^{(i)}$ are well-defined and satisfy $v_{\mu}\geq v_{\lambda}$ whenever $\emptyset\subset\mu\subset\lambda\subseteq[n]\setminus\{i\}$. Thus, Lemma C.2 applies once more, meaning the events $\{B_{\mu}(\mathbf{v}^{(i)}):\mu\subseteq[n]\setminus\{i\}\}$ partition the sample space $\Omega$, and we have
$\sum_{\mu\subseteq[n]\setminus\{i\}}\mathbb{P}(B_{\mu}(\mathbf{v}^{(i)}))=1.$
This simplifies further the inequality to
$\mathbb{P}\Bigg{(}\sum_{i\in[n]}\theta_{i}X_{i}>x\Bigg{)}\geq\sum_{i=1}^{n}\theta_{i}\overline{F}_{i}(x),$
which completes the proof.
Acknowledgements
The author thanks Hansjörg Albrecher for helpful comments on an earlier version of this manuscript.